OpenStudy (anonymous):

It takes you 5 min to walk with an average velocity of .75 m/s to the north from the parking lot to the entrance of the amusement park. What is your displacement? a. 400 m to the north b. 3.75 m to the east c. 225 m to the north d. need more information

OpenStudy (anonymous):

@ooops @Hero @Preetha @paki

OpenStudy (paki):

have you any idea...?

OpenStudy (anonymous):

Nope, still nothing.. :/

OpenStudy (paki):

Velocity = (Displacement) / (Time) of D=V/T.... when we put values here... the required answer is not obtained...

OpenStudy (anonymous):

You have the time it takes, 5 Minutes, to walk from A to B with the average velocity given as .75 m /s. From the formula of Velocity you can solve for the displacement during that time interval. \[D = V \times T\] Where D = Distance in meters V = Velocity in meters/sec T = Time in seconds Seeing as your time is in minutes you must first convert to seconds. We can use conversion factors for that: \[minutes \times \frac{ seconds }{ minutes } = seconds\] Placing the proper Minutes and Seconds into your conversion factor you will solve for seconds and then you can substitute that back into your equation\[D = V \times T\] @teddyshakur